1.What is the total GDP for South Africa in rands assuming that there are people in South Africa that the GDP CAPITA is $11900 and the rand dollar exchange rate is R6.50 to the dollar.
2. If you invest R10 000 at 10% interest now ,how much will it be worth after 20 years?
Thank you in advance.

Answer by Boreal(15235)

The first needs the population of RSA. 11,900*6.5 R/$=77,350 R. per capita.
The total GDP for RSA in Rands is 77,350*number of people.
--------------------
Assuming one compounding per year,
P=10,000(1+0.1)^20; 1.1^20=6.73
without rounding multiply by 10000=67,275 rands, round at the end.